An individual can win a bouncy ball by guessing under which one of four cups the ball is located. After each guess, if the ball is won, a new ball is placed randomly under one of the four cups. If the ball is not won, then the ball is again placed randomly under one of the four cups. If an individual makes five guesses, what is the probability the individual will win a prize exactly two times?
Answer #1

This will be a binomial distribution with parameter:

n = 5, p = P(Winning) = 1/4 = 0.25

Hence,

P(Exactly two times)

= P(X = 2)

= 0.2637